DQRID : D990906.1
Start DateStart TimeEnd DateEnd Time Data Quality Metric
04/15/1999000009/02/19992221Does not affect quality
SGP/MFR10M/C1 - Head heater not working
DataStreams:sgpmfr10mC1.a0, sgpmfr10mC1.b1
On 19990415, the tower upon which the 10-m MFR was mounted was dismantled and replaced 
with a so-called "tippy tower".  After the 10-m MFR was reinstalled on the new tower, the 
mfr head heater did not work, or worked intermittently (at best).  A new mfr head was 
installed on 99.09.02 and that solved the problem.  

Data was collected from the old head in the morning up to the time it was replaced.  Once 
the new head was installed, data collection resumed.  There's a gap of about an hour 
between the time the old head was removed and the new head was installed.  If we try to gauge 
how well the old head was calibrated by comparing the two segments of data (old head, 
new head) it appears as if the old head remained in a good state of calibration even though 
the head heater was not working.
  • head_temp
  • head_temp

Close this window